OpenStudy (anonymous):

A series circuit consists of a 12.0 V source of emf, a 2.00 mF capacitor, a 1000 Ω resistor, and a switch. When the switch is closed, how long does it take for the current to reach one-tenth its maximum value?

OpenStudy (anonymous):

(a) 4.61 s (b) 2.30 s (c) 18.0 s (d) 9.00 s (e) 0.693 s

OpenStudy (anonymous):

See this and try by yourself http://hyperphysics.phy-astr.gsu.edu/hbase/electric/capchg.html

OpenStudy (anonymous):

@shivam_bhalla You can answer it?

OpenStudy (anonymous):

See the formula is \[\huge {I = {{v_0} \over {t}}{e^{{-t}/{RC}}}}\] You know what is V, R ,C , I from the question Now substitute and find t . If you have doubt in simplification, then post your step one by one and I will help you

OpenStudy (anonymous):

So i will then substitute for V_0 with 12 and C with 20 mF and R with 1000 ohms, what about the current and when I simplify the equation, it will be like this : t = V_0 / I * e^-t/RC, so how can i substitute for the power -t and is e for exponential? Thanks

OpenStudy (anonymous):

Yes :) Make sure you write C=20 mF = 20 * 10^(-3) F

OpenStudy (anonymous):

And what should i substitute for I?

OpenStudy (anonymous):

\[I = { 1 \over 10} {v_b \over R}\]

OpenStudy (anonymous):

Since the maximimum current in the circuit is \[I_{\max} = {v_b \over R}\] and it is given in the question that I = 1/10 of I_max

OpenStudy (anonymous):

Yes you are right :). Well now the power (-t) for the exponential is the problem because that what I need indeed :) so what should I do with?

OpenStudy (anonymous):

The best way is to take log_e on both sides and then apply power rule. Power rule is \[\large \log_a b^c = c*\log_a b\]

OpenStudy (anonymous):

@benanderson5900 , you got the answer because I am getting the answer :)

OpenStudy (anonymous):

Hahaha well I just tried but I got irrelevant numbers to the onesI have to choose from :D

OpenStudy (anonymous):

I am getting one of your choices, show me your working out so that I can help you out. I am getting option A

OpenStudy (anonymous):

Well I think the problem is still in log. After taking log both sides, how the equation should be looks like?

OpenStudy (anonymous):

Sorry it should be \[\large \log_e {1 \over 10} = \log_e e^{-t/2}\]

OpenStudy (anonymous):

Now apply power rule and tell me your equation. (For power rule see above in this thread)

OpenStudy (anonymous):

Well so log will cancel the exponential, so the equation will be "log(1/10) = -t/2?

OpenStudy (anonymous):

Exactly. Then One more rule which is \[\large \log_a {p \over q} = \log_a p - \log_a q\]

OpenStudy (anonymous):

Well so log(1/10) = log(1) - log(10)?

OpenStudy (anonymous):

yes and what is log(1) =??

OpenStudy (anonymous):

0

OpenStudy (anonymous):

Yes now tell me your final equation after doing all this.

OpenStudy (anonymous):

-1 = -t/2?

OpenStudy (anonymous):

No. \[\large \log_e 10 \neq 1 \]

OpenStudy (anonymous):

Sorry, equals to 2.303?

OpenStudy (anonymous):

Yes :) Now you will get your required answer :)

OpenStudy (anonymous):

Haha well yeah I just got it :). Thanks a lot really :)

OpenStudy (anonymous):

Anytime :D

OpenStudy (anonymous):

Actually it is no the best answer but the best help and support instead :D and that what matters :).

OpenStudy (anonymous):

Exactly :D. I will give you a medal for that :)

OpenStudy (anonymous):